Age of Suspect

11

18

18

13

22

22

21

30

24

Age of Victim

13

14

18

15

22

22

29

44

29

Age of Suspect

45

27

59

33

49

51

31

64

23

Age of Victim

36

37

42

43

47

51

49

61

51

Age of Suspect and Age of Victim in Single Suspect Incidences of Gun Violence

The table above shows the age of the suspect and the age of the victim in a random sample of 18 gun violence incidents that occurred in Utah between January of 2013 and March of 2018.

1. Based on the variables involved in this relationship which variable do you think is the explanatory (x) variable and which is the response (y) variable?

2. Calculate the correlation between the two variables. r=

3. Interpret the full meaning of the correlation coefficient you calculated in #2, including direction, strength, and relationship between variables.

4. Calculate the average and SD for the variable you chose as the explanatory variable.

Average =

SD =

5. Calculate the average and SD for the variable you chose as the response variable.

Average=

SD=

6. Find the equation of the regression line that fits your data. Show all calculation.

7. Interpret the meaning of the slope of your regression model from question #6

8. Interpret the meaning of the y-intercept of your regression model from question #6. If there is no practical meaning, explain why.

9. Demonstrate how someone might use the regression model you found in question #6 to predict the value of a response variable. That is, plug a hypothetical x-value in your model and explain what it predicts.
Answer #1

( 1 ) Explanatory (x) variable is = Age of Suspect

response (y) variable is =  Age of Victim

( 2 )

Find X⋅Y , X2 and Y2 as it was done in the table below.

X Y X⋅Y X⋅X Y⋅Y  
11 13 143 121 169
18 14 252 324 196
18 18 324 324 324
13 15 195 169 225
22 22 484 484 484
22 22 484 484 484
21 29 609 441 841
30 44 1320 900 1936
24 29 696 576 841
45 36 1620 2025 1296
27 37 999 729 1369
59 42 2478 3481 1764
33 43 1419 1089 1849
49 47 2303 2401 2209
51 51 2601 2601 2601
31 49 1519 961 2401
64 61 3904 4096 3721
23 51 1173 529 2601

Find the sum of every column to get:

ΣΧ = 561, ΣΥ= 623, ΣΧ.Υ = 22523, Σχ? – 21735, ΣΥ? – 25311 Use the following formula to work out the correlation coefficient.

( 3 ) There is strong positive linear relationship between variables.

( 4 )

explanatory variable.

Average = 561 / 18 =  31.1667

SD = 15.8123

Σ(αι - Χ)* 4250.5 4250.5 18 - 15.8123

( 5 )

response variable.

Average = 623 / 18 = 34.6111

SD = 14.8488

Σ(1, - ) Ν η 1 3748.2778 9 8 14.8488 18 - 1

( 6 )

Find the sum of every column Σx = 561, ΣΥ = 623, ΣΧ. Υ = 22523, Σx2 - 21735 Use the following equations to find a and b. . ΣΥ